# Alignment-Sensitive Minimax Rates for Spectral Algorithms with Learned Kernels

### v2

@@ 10 Discussion
Quantifying adaptivity. Like the sparsity level in high-dimensional statistics, the ESD is a population quantity for theoretical analyses rather than an input to training. It serves as a quantitative target for adaptive algorithms on the population level: by comparing the ESD of a particular signal w.r.t. different kernels, we can determine which kernel permits better generalization for this signal.
Connecting adaptivity and generalization. Our span profile framework clarifies why adaptive machine learning methods often outperform classical fixed-kernel approaches. In classical methods with a fixed kernel spectrum $\bm{\lambda}^{(0)}$, the target signal $\bm{\theta}^{*}$ might exhibit poor alignment, resulting in a large span profile $D_{\bm{\theta}^{*},\bm{\lambda}^{(0)}}$. Consequently, the signal resides in a class $\mathcal{F}_{\mathbf{K}^{(0)},\bm{\lambda}^{(0)}}$ with a large quota sequence $\mathbf{K}^{(0)}$, which implies high minimax risk. By contrast, adaptive methods modify the kernel during training. Successful adaptation improves the alignment by reducing the span profile $D_{\bm{\theta}^{*},\bm{\lambda}^{(a)}}$ of the same signal w.r.t. the adapted kernel spectrum $\bm{\lambda}^{(a)}$. This adaptation places the signal in a class $\mathcal{F}_{\mathbf{K}^{(a)},\bm{\lambda}^{(a)}}$ with a smaller quota sequence $\mathbf{K}^{(a)}$, which implies lower minimax risk.
@@ A.2 Proof on Minimax Results
For any $v\in\mathcal{V}$, the parameter vector $\bm{\theta}^{(v)}$ constructed in Equation˜32 lies in $\mathcal{F}_{\bm{K},\bm{\lambda}}$.
@@
Let $\delta=\sqrt{c\sigma_{0}^{2}/n}$ with the constant $c=\tfrac{1}{4}$. Let $B_{n}=\{\pi_{1},\dots,\pi_{K_{n}}\}$. We define the collection of hypercubes vertices $\mathcal{V}=\{-1,1\}^{K}$. For every vertex $v\in\mathcal{V}$, define a parameter vector $\bm{\theta}^{(v)}=(\theta_{j}^{(v)})_{j=1}^{d}$ as in Equation˜32. There are $2^{K}$ such vectors $\{\bm{\theta}^{(v)}\}$. For each $v\in\mathcal{V}$, let $P_{v}$ be the sampling distribution of the sequence model in Equation˜4 with $\bm{\theta}^{*}=\bm{\theta}^{(v)}$, $\sigma^{2}=\sigma_{0}^{2}/n$, and $\{\xi_{j}\}_{j\in[d]}$ being i.i.d. normal. Let $\rho$ be Hamming distance on $\mathcal{V}$. The rest of the proof is identical to that of Theorem˜4.3 and is omitted. ∎
A.3 Details of Examples in Equation˜9
We provide the details of Equation˜9 for illustration of the concepts of ESD and span profile through several examples.
@@ A.3 Details of Examples in Equation˜9
For any $n\geq 3$, let $k=K_{n}$. By definition of the ceiling function, $k\geq(\log n)^{1/b}$, which implies $k^{b}\geq\log n$, and thus $e^{k^{b}}\geq n$. Then, $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)=\sigma_{0}^{2}e^{-k^{b}}\leq\sigma_{0}^{2}/n$. By Proposition˜3.5, we have $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma_{0}^{2}/n)\leq k=K_{n}$.
@@
Let’s assume a polynomial eigenvalue decay $\lambda_{j}\asymp j^{-\gamma}$ for some $\gamma>0$. Let $S$ be the left hand side of Equation˜34. Since $\theta_{1}^{*}=0$, we have
@@
Appendix B Proofs for results in Section˜8
Proof of Proposition˜8.3.
@@ Appendix B Proofs for results in Section˜8
Proof of Theorem˜8.5.
Upper bound: The upper bound follows the proof of the upper bound in Proposition˜8.3. To see this, we note that since $f^{*}\in\mathcal{F}_{K,\bm{\lambda},n}$, we have $\|f\|_{\infty}^{2}\leq\sigma_{0}^{2}C_{0}^{2}$. Therefore, $\sigma^{2}\leq\bar{\sigma}^{2}/n$. We can then apply the argument in Proposition˜8.3 with $\sigma^{2}$ replaced by $\bar{\sigma}^{2}/n$.
@@
For each $v\in\mathcal{V}$, let $P_{v}$ be the sampling distribution of $\{z_{i}=(x_{i},y_{i})\}_{i\leq n}$ from the regression model Equation˜20 with $f^{*}=f^{(v)}$. Let $\rho$ be the Hamming distance on $\mathcal{V}$. If $v$ and $w\in\mathcal{V}$ differ in exactly one coordinate (i.e., $\rho(v,w)=1$), then
@@
the upper bound $\bar{\sigma}^{2}K_{n}/n$ is immediately implied by Theorem˜8.5.
The lower bound follows the same argument as in the proof of Theorem˜4.3, but replace the construction of parameter vectors in Equation˜32 by the construction of functions in Equation˜35. Following the proof for Theorem˜4.3, we use ˜4.1 to ensure the constructed functions all belong to $\mathcal{F}_{\mathbf{K},\mathbf{k}}$. Then the lower bound is given using Assouad’s Lemma as in the proof of Theorem˜8.5. Below, we provide the details for completeness.
@@
Fix $n$ and set $m:=\lfloor c_{1}K_{n}\rfloor$ where $c_{1}$ comes from ˜8.4.
@@
Since $\mathbf{k}$ is $(K_{n},n)$-regular, Equation˜36 holds and reads as
@@
If $\gamma_{2}C_{1}n\kappa^{2}\leq\sigma_{0}^{2}C_{0}^{2}$, then $\|f^{(v)}\|_{\infty}^{2}\leq\sigma_{0}^{2}C_{0}^{2}$. Furthermore, if $m\gamma^{2}\leq(2n)^{-1}\sigma_{0}^{2}K_{n}$, we can the same argument in Lemma˜A.2 (in particular, using ˜4.1 to dervie Equation˜33) to show that $f^{(v)}\in\mathcal{F}_{\mathbf{K},\mathbf{k}}$.
@@
We then follows the same argument in the proof of lower bound in Theorem˜8.5 to obtain
@@
Appendix C Proofs for Result on overparameterized gradient flow
@@
where $\tilde{\sigma}^{2}\leq\min\{\frac{|B_{1}|}{C_{B_{1}}}\varepsilon^{2},c^{\prime}\varepsilon^{2}\}$, $C_{B_{1}}=\min\{1\vee(|A_{1}|-|B_{1}|),|B_{2}|\}$ and $c^{\prime}$ is a constant $\leq 1$. Also recall from the assumption of Theorem˜5.2 that suppose $|\theta^{*}_{i}|>|\theta^{*}_{j}|>M$, then if $C_{max}M>|\theta^{*}_{i}|$, $\eta_{i,j}=|\theta_{i}^{*}|-|\theta_{j}^{*}|>C_{\eta}\varepsilon^{\prime}$; otherwise, $\frac{|\theta_{i}^{*}|}{|\theta_{j}^{*}|}>(1+\frac{c_{\eta}}{D})$.
@@
Part 1: Very small eigenvalues can be ignored. From ˜5.1, we have $\inf_{j\in S}\lambda_{j}>n^{-\delta}$, where $\delta$ is a constant. For $i\in R$, if $\lambda_{i}<n^{-(2.1\delta\vee 5)}$, Proposition˜C.2 implies that at $t_{2}$, we have
@@
If $j\in B_{2}$, by Proposition C.1, we have $\left|\theta_{j}(t_{2})-\theta_{j}^{*}\right|<\varepsilon^{\prime}\leq\frac{1}{C_{M}}M$. We apply Equation˜50 to get $\beta^{2}=a^{2}-\lambda<a^{2}$ and $\beta^{2}=D^{-1}(b^{2}-b_{0}^{2})<D^{-1}b^{2}$, and thus
@@
It follows that $\tilde{\lambda}_{j}(t_{2})\geq cD^{\frac{D}{(D+2)}}M^{\frac{2(D+1)}{D+2}}$. Moreover, for $\lambda_{j}$ and $D^{-1}b_{0}^{2}$ that are much smaller than $c\cdot D^{-\frac{D}{D+2}}\cdot M^{\frac{2}{D+2}}$, we use Equation˜50 to obtain
@@
It follows that $a_{i}^{2}(t_{2})b_{i}^{2D}(t_{2})>a_{j}^{2}(t_{2})b_{j}^{2D}(t_{2})$. Combined with Equation˜50, we have
@@
Recall that $|\beta_{i}(t_{2})|<\frac{1}{C_{D}}|\beta_{j}(t_{2})|$ and $\beta_{j}^{2}(t_{2})>CD^{-\frac{2}{D+2}}M^{\frac{2}{D+2}}$. If we choose the constant $C_{D}$ such that $(1+c(C_{D}-1))^{D}$ is large enough, the inequality Equation˜40 will implies $\lambda_{i}$ larger than its upper bound in the definition of set $A_{1}$. (We can let $C_{D}=1+c\cdot\frac{1}{D}$.)
@@
If $\lambda_{i}\geq\lambda_{j}$, by Equation˜38 and monotonicity, we always have $\tilde{\lambda}_{i}(t)>\tilde{\lambda}_{j}(t)$.
@@
If $\tilde{\lambda}_{i}(t)\leq\tilde{\lambda}_{j}(t)$, then $\beta_{i}(t_{2})>C_{D}\beta_{j}(t_{2})$. By Equation˜39, both $\beta_{i}(t_{2})$ and $\beta_{j}(t_{2})$ are much larger then $D^{-\frac{1}{2}}b_{0}$. It then follows that
@@
where $C^{*}$ is a constant that can be made large enough by choosing $C_{D}$. Therefore, using the same reason as in Equation˜40, we can get
@@
In the following, let $j\in B_{1}$. By $y=\theta^{*}+\xi$, we have $|z_{j}|<\tilde{\sigma}+\varepsilon^{\prime}$. By Equation˜53 and $|\theta_{j}|<|z_{j}|$, we have
@@
    W.L.O.G., we can write divide $A_{11}$ into two subsets such that $|A_{111}|=|B_{2}|$ and $|A_{112}|=|B_{11}|$. We can then write Equation˜41 as
@@
    Note the assumption of Theorem˜5.2 that $(|B_{2}|+C_{B_{1}})\tilde{\sigma}^{2}\leq|B_{2}|M^{2}$. We add the last two inequalities Equations˜43 and 43 together to get
@@
    Note that at time $t_{1}$, the elements in the later part are in $B_{1}$, $B_{21}$, and $B_{22}:=B_{2}\setminus B_{21}$. Therefore, Equation˜41 can be written as
@@
    By definition of $B_{2}$, $A_{1}$, and $A_{3}$, each squared element in $B_{2}$ is larger than that of both $A_{1}$ and $A_{3}$. In addition, since $A_{2}$ and $B_{2}$ will be monotonously nonincreasing, for any element in $B_{22}$, its squared signal will be no less than that of any element in $\theta_{A_{21}}$. Therefore, we conclude that Equation˜45 is nonnegative and will be positive if $B_{22}$ is not empty.

### v3

@@ 11 Discussion
Like the sparsity level in high dimensional statistics, ESD is a population quantity for theory rather than a training input. It is useful as a quantitative diagnostic: comparing ESD across kernels clarifies which representation supports better generalization for a given signal.
@@ A.2 Proof on Minimax Results
For any $v\in\mathcal{V}$, the parameter vector $\bm{\theta}^{(v)}$ constructed in Equation˜33 lies in $\mathcal{F}_{\bm{K},\bm{\lambda}}$.
@@
Let $\delta=\sqrt{c\sigma_{0}^{2}/n}$ with the constant $c=\tfrac{1}{4}$. Let $B_{n}=\{\pi_{1},\dots,\pi_{K_{n}}\}$. We define the collection of hypercubes vertices $\mathcal{V}=\{-1,1\}^{K}$. For every vertex $v\in\mathcal{V}$, define a parameter vector $\bm{\theta}^{(v)}=(\theta_{j}^{(v)})_{j=1}^{d}$ as in Equation˜33. There are $2^{K}$ such vectors $\{\bm{\theta}^{(v)}\}$. For each $v\in\mathcal{V}$, let $P_{v}$ be the sampling distribution of the sequence model in Equation˜4 with $\bm{\theta}^{*}=\bm{\theta}^{(v)}$, $\sigma^{2}=\sigma_{0}^{2}/n$, and $\{\xi_{j}\}_{j\in[d]}$ being i.i.d. normal. Let $\rho$ be Hamming distance on $\mathcal{V}$. The rest of the proof is identical to that of Theorem˜4.3 and is omitted. ∎
A.3 Details of Example˜3.4
We provide the details of Example˜3.4 for illustration of the concepts of ESD and span profile through several examples.
@@
For any $n\geq 3$, let $k=K_{n}$. By definition of the ceiling function, $k\geq(\log n)^{1/b}$, which implies $k^{b}\geq\log n$, and thus $e^{k^{b}}\geq n$. Then, $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)=\sigma_{0}^{2}e^{-k^{b}}\leq\sigma_{0}^{2}/n$. By Proposition˜3.6, we have $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma_{0}^{2}/n)\leq k=K_{n}$.
@@
Let’s assume a polynomial eigenvalue decay $\lambda_{j}\asymp j^{-\gamma}$ for some $\gamma>0$. Let $S$ be the left hand side of Equation˜35. Since $\theta_{1}^{*}=0$, we have
@@
Appendix B Proofs for results in Section˜9
Proof of Proposition˜9.3.
@@ Appendix B Proofs for results in Section˜9
Proof of Theorem˜9.5.
Upper bound: The upper bound follows the proof of the upper bound in Proposition˜9.3. To see this, we note that since $f^{*}\in\mathcal{F}_{K,\bm{\lambda},n}$, we have $\|f\|_{\infty}^{2}\leq\sigma_{0}^{2}C_{0}^{2}$. Therefore, $\sigma^{2}\leq\bar{\sigma}^{2}/n$. We can then apply the argument in Proposition˜9.3 with $\sigma^{2}$ replaced by $\bar{\sigma}^{2}/n$.
@@
For each $v\in\mathcal{V}$, let $P_{v}$ be the sampling distribution of $\{z_{i}=(x_{i},y_{i})\}_{i\leq n}$ from the regression model Equation˜21 with $f^{*}=f^{(v)}$. Let $\rho$ be the Hamming distance on $\mathcal{V}$. If $v$ and $w\in\mathcal{V}$ differ in exactly one coordinate (i.e., $\rho(v,w)=1$), then
@@
the upper bound $\bar{\sigma}^{2}K_{n}/n$ is immediately implied by Theorem˜9.5.
The lower bound follows the same argument as in the proof of Theorem˜4.3, but replace the construction of parameter vectors in Equation˜33 by the construction of functions in Equation˜36. Following the proof for Theorem˜4.3, we use ˜4.1 to ensure the constructed functions all belong to $\mathcal{F}_{\mathbf{K},\mathbf{k}}$. Then the lower bound is given using Assouad’s Lemma as in the proof of Theorem˜9.5. Below, we provide the details for completeness.
@@
Fix $n$ and set $m:=\lfloor c_{1}K_{n}\rfloor$ where $c_{1}$ comes from ˜9.4.
@@
Since $\mathbf{k}$ is $(K_{n},n)$-regular, Equation˜37 holds and reads as
@@
If $\gamma_{2}C_{1}n\kappa^{2}\leq\sigma_{0}^{2}C_{0}^{2}$, then $\|f^{(v)}\|_{\infty}^{2}\leq\sigma_{0}^{2}C_{0}^{2}$. Furthermore, if $m\gamma^{2}\leq(2n)^{-1}\sigma_{0}^{2}K_{n}$, we can the same argument in Lemma˜A.2 (in particular, using ˜4.1 to dervie Equation˜34) to show that $f^{(v)}\in\mathcal{F}_{\mathbf{K},\mathbf{k}}$.
@@
We then follows the same argument in the proof of lower bound in Theorem˜9.5 to obtain
@@
B.1 Proofs for results in Section˜5
Proof of Proposition˜5.1.
We note that for the top $k$ terms, $\lambda_{(i)}\geq\nu$, so $\frac{\lambda_{(i)}}{\lambda_{(i)}+\nu}\geq\frac{1}{2}$ and hence
$\sum_{i=1}^{k}\Bigl(\frac{\lambda_{(i)}}{\lambda_{(i)}+\nu}\Bigr)^{2}\geq\frac{k}{4}.$
For the tail terms, $\lambda_{(i)}<\nu$ for $i\geq k+1$, so $\frac{\lambda_{(i)}}{\lambda_{(i)}+\nu}\geq\frac{\lambda_{(i)}}{2\nu}$ and thus
$\sum_{i=k+1}^{d}\Bigl(\frac{\lambda_{(i)}}{\lambda_{(i)}+\nu}\Bigr)^{2}\geq\frac{1}{4\nu^{2}}\sum_{i=k+1}^{d}\lambda_{(i)}^{2}.$
Combining the two cases, we have
$V_{\mathrm{R}}(\nu)$ $=\sigma^{2}\sum_{i=1}^{d}\Bigl(\frac{\lambda_{(i)}}{\lambda_{(i)}+\nu}\Bigr)^{2}$
$\geq\sigma^{2}\Biggl(\frac{k}{4}+\frac{1}{4v^{2}}\sum_{i=k+1}^{d}\lambda_{(i)}^{2}\Biggr)$
$\geq\frac{1}{4}\sigma^{2}\Biggl(k+\sum_{i=k+1}^{d}\Bigl(\frac{\lambda_{(i)}}{\lambda_{(k)}}\Bigr)^{2}\Biggr).$
Proof of Proposition˜5.3.
For any $\nu\leq\lambda_{(i_{0})}$, we have $\lambda_{(i)}+\nu\leq 2\lambda_{(i)}$ for all $i\leq i_{0}$. We then have
$B_{\mathrm{R}}(\nu)\geq\sum_{i=1}^{i_{0}}\Bigl(\frac{\nu}{\lambda_{(i)}+\nu}\Bigr)^{2}(\theta^{*}_{\pi_{i}})^{2}\geq\frac{\nu^{2}}{4}\sum_{i=1}^{i_{0}}\frac{(\theta^{*}_{\pi_{i}})^{2}}{\lambda_{(i)}^{2}}\geq c_{r}\,\nu^{2},$
where the last inequality is due to ˜5.2. ∎
Proof of Theorem˜5.6.
Fix $\nu\in(0,\lambda_{(i_{0})}]$ and let $k:=k_{\bm{\lambda}}(\nu)$. By definition of $k_{\bm{\lambda}}(\nu)$, we have $\lambda_{(k)}\geq\nu>\lambda_{(k+1)}$ and, since $\nu\leq\lambda_{(i_{0})}$, we have $k\geq i_{0}$.
Under ˜5.2, for any $\nu\leq\lambda_{(i_{0})}$ we already showed in Equation˜12 that
$B_{\mathrm{R}}(\nu)\geq c_{r}\nu^{2}.$
Proposition˜5.1 implies that
$V_{\mathrm{R}}(\nu)\geq\frac{1}{4}\widetilde{V}(k).$
Combining the above two lower bounds, we have
$\mathcal{R}_{\mathrm{R}}(\nu)=B_{\mathrm{R}}(\nu)+V_{\mathrm{R}}(\nu)\geq\max\Bigl\{B_{\mathrm{R}}(\nu),V_{\mathrm{R}}(\nu)\Bigr\}\geq\max\Bigl\{c_{r}\nu^{2},\frac{1}{4}\widetilde{V}(k)\Bigr\}.$
Recall
$d^{\Delta}=\max\Bigl\{m:\ c_{r}\,\lambda_{(m)}^{2}>\sigma^{2}\,m\bigl(1+\widetilde{N}(m,d;\bm{\lambda})\bigr)\Bigr\}=\max\Bigl\{m:\ c_{r}\,\lambda_{(m)}^{2}>\widetilde{V}(m)\Bigr\}.$
We consider two cases.
Case 1: $k\geq d^{\Delta}$. Then $\widetilde{V}(k)\geq\sigma^{2}k\geq\sigma^{2}d^{\Delta}$, so
$\mathcal{R}_{\mathrm{R}}(\nu)\geq\frac{1}{4}\widetilde{V}(k)\geq\frac{1}{4}\sigma^{2}d^{\Delta}.$
Case 2: $k\leq d^{\Delta}-1$. Since $\nu>\lambda_{(k+1)}$ and $k+1\leq d^{\Delta}$, monotonicity of the spectrum gives $\lambda_{(k+1)}\geq\lambda_{(d^{\Delta})}$, hence $\nu\geq\lambda_{(d^{\Delta})}$ and
$\mathcal{R}_{\mathrm{R}}(\nu)\geq c_{r}\nu^{2}\geq c_{r}\lambda_{(d^{\Delta})}^{2}.$
By definition of $d^{\Delta}$, we have $c_{r}\lambda_{(d^{\Delta})}^{2}>\widetilde{V}(d^{\Delta})\geq\sigma^{2}d^{\Delta}$, so
$\mathcal{R}_{\mathrm{R}}(\nu)\geq c_{r}\lambda_{(d^{\Delta})}^{2}>\sigma^{2}d^{\Delta}\geq\frac{1}{4}\sigma^{2}d^{\Delta}.$
In both cases, $\mathcal{R}_{\mathrm{R}}(\nu)\geq\frac{1}{4}\sigma^{2}d^{\Delta}$. Taking the infimum over $\nu\in(0,\lambda_{(i_{0})}]$ yields the second display. ∎
Appendix C Proofs for results on overparameterized gradient flow
@@
where $\tilde{\sigma}^{2}\leq\min\{\frac{|B_{1}|}{C_{B_{1}}}\varepsilon^{2},c^{\prime}\varepsilon^{2}\}$, $C_{B_{1}}=\min\{1\vee(|A_{1}|-|B_{1}|),|B_{2}|\}$ and $c^{\prime}$ is a constant $\leq 1$. Also recall from the assumption of Theorem˜6.2 that suppose $|\theta^{*}_{i}|>|\theta^{*}_{j}|>M$, then if $C_{max}M>|\theta^{*}_{i}|$, $\eta_{i,j}=|\theta_{i}^{*}|-|\theta_{j}^{*}|>C_{\eta}\varepsilon^{\prime}$; otherwise, $\frac{|\theta_{i}^{*}|}{|\theta_{j}^{*}|}>(1+\frac{c_{\eta}}{D})$.
@@
Part 1: Very small eigenvalues can be ignored. From ˜6.1, we have $\inf_{j\in S}\lambda_{j}>n^{-\delta}$, where $\delta$ is a constant. For $i\in R$, if $\lambda_{i}<n^{-(2.1\delta\vee 5)}$, Proposition˜C.2 implies that at $t_{2}$, we have
@@
If $j\in B_{2}$, by Proposition C.1, we have $\left|\theta_{j}(t_{2})-\theta_{j}^{*}\right|<\varepsilon^{\prime}\leq\frac{1}{C_{M}}M$. We apply Equation˜51 to get $\beta^{2}=a^{2}-\lambda<a^{2}$ and $\beta^{2}=D^{-1}(b^{2}-b_{0}^{2})<D^{-1}b^{2}$, and thus
@@
It follows that $\tilde{\lambda}_{j}(t_{2})\geq cD^{\frac{D}{(D+2)}}M^{\frac{2(D+1)}{D+2}}$. Moreover, for $\lambda_{j}$ and $D^{-1}b_{0}^{2}$ that are much smaller than $c\cdot D^{-\frac{D}{D+2}}\cdot M^{\frac{2}{D+2}}$, we use Equation˜51 to obtain
@@
It follows that $a_{i}^{2}(t_{2})b_{i}^{2D}(t_{2})>a_{j}^{2}(t_{2})b_{j}^{2D}(t_{2})$. Combined with Equation˜51, we have
@@
Recall that $|\beta_{i}(t_{2})|<\frac{1}{C_{D}}|\beta_{j}(t_{2})|$ and $\beta_{j}^{2}(t_{2})>CD^{-\frac{2}{D+2}}M^{\frac{2}{D+2}}$. If we choose the constant $C_{D}$ such that $(1+c(C_{D}-1))^{D}$ is large enough, the inequality Equation˜41 will implies $\lambda_{i}$ larger than its upper bound in the definition of set $A_{1}$. (We can let $C_{D}=1+c\cdot\frac{1}{D}$.)
@@
If $\lambda_{i}\geq\lambda_{j}$, by Equation˜39 and monotonicity, we always have $\tilde{\lambda}_{i}(t)>\tilde{\lambda}_{j}(t)$.
@@
If $\tilde{\lambda}_{i}(t)\leq\tilde{\lambda}_{j}(t)$, then $\beta_{i}(t_{2})>C_{D}\beta_{j}(t_{2})$. By Equation˜40, both $\beta_{i}(t_{2})$ and $\beta_{j}(t_{2})$ are much larger then $D^{-\frac{1}{2}}b_{0}$. It then follows that
@@
where $C^{*}$ is a constant that can be made large enough by choosing $C_{D}$. Therefore, using the same reason as in Equation˜41, we can get
@@
In the following, let $j\in B_{1}$. By $y=\theta^{*}+\xi$, we have $|z_{j}|<\tilde{\sigma}+\varepsilon^{\prime}$. By Equation˜54 and $|\theta_{j}|<|z_{j}|$, we have
@@
    W.L.O.G., we can write divide $A_{11}$ into two subsets such that $|A_{111}|=|B_{2}|$ and $|A_{112}|=|B_{11}|$. We can then write Equation˜42 as
@@
    Note the assumption of Theorem˜6.2 that $(|B_{2}|+C_{B_{1}})\tilde{\sigma}^{2}\leq|B_{2}|M^{2}$. We add the last two inequalities Equations˜44 and 44 together to get
@@
    Note that at time $t_{1}$, the elements in the later part are in $B_{1}$, $B_{21}$, and $B_{22}:=B_{2}\setminus B_{21}$. Therefore, Equation˜42 can be written as
@@
    By definition of $B_{2}$, $A_{1}$, and $A_{3}$, each squared element in $B_{2}$ is larger than that of both $A_{1}$ and $A_{3}$. In addition, since $A_{2}$ and $B_{2}$ will be monotonously nonincreasing, for any element in $B_{22}$, its squared signal will be no less than that of any element in $\theta_{A_{21}}$. Therefore, we conclude that Equation˜46 is nonnegative and will be positive if $B_{22}$ is not empty.